Solve
8x-(3x+6)=29
X=?
I know this is simple but I'm not sure my answer is correct, can someone help? thank you

Found 2 solutions by richwmiller, k3v1nl4m:
Answer by richwmiller(17219)

First you should tell us what you have tried.
Second have you even looked at other problems in this category to see if you could learn something.
Remove the parentheses
we must remember
a-minus times a plus = minus
b-if there is no sign it is plus
c-minus times a minus =plus
Example: -5*+4=-20
That could be written in other ways.
-5(4)=-20
4(-5)=-20
-5(+4)=-20

Answer by k3v1nl4m(1)

8x-(3x+6)=29
8x-3x-6=29 (distributed negative sign)
5x-6=29 (simplify x)
5x-6+6=29+6 (add 6 to both sides)
5x=35 (divide by 5)
x=7